OpenStudy (anonymous):

One of two urns is chosen at random with one as likely to be chosen as other. Then a ball is drawn from the ch? one of two urns is chosen at random with one as likely to be chosen as other. Then a ball is drawn from the chosen urn. Urn 1 contains 1 white and 4 red balls. and urn 2 has 4 white and 1 red ball. If a white ball is drawn, what is the probability it came from urn 2?

OpenStudy (anonymous):

I might be a bit rusty, but I believe the answer is 4/10

OpenStudy (anonymous):

you are given that a while ball is drawn, no probability is to be calculated here. there are 5 white balls, 4 of which are in urn 2. the probability is 4/5, not 4/10. typo on my part

OpenStudy (anonymous):

1/2 * 4/5= 4/10 =0.4 It is the probability of choosing urn 2 multiply the probability of getting a white ball.

OpenStudy (swissgirl):

Do you follow what ghost answered?